# SReferences

### v2

@@ I Introduction
Here, we present a fusion-based architecture tailored to quantum emitters and analyze how relevant physical processes affect its fault-tolerant performance. The architecture is based on the foliated Floquet color code (FFCC) [11] where each resource state can be deterministically generated by a single quantum emitter, with high tolerance to photon loss and qubit errors. We develop a framework to describe how the dominant noise processes in quantum emitters propagate to the logical qubit level of the architecture and analyze its performance in terms of fault-tolerant thresholds to physical noises. Specifically, we investigate photon distinguishability, Markovian spin noise in spin-photon interfaces, and photon loss. We show that our tailored architectures are highly resilient to these noises, as summarized in Table 1.
@@
Table 1: Summary of main results. sFFCC: synchronous foliated Floquet color code lattice. REP: logical fusion with repetition code of fixed code size $m$. RUS: repeat-until-success logical fusion with varying code size $m\leq N$ where $N$ is the maximum number of physical fusion attempts. $\ell$, $V$ and $p^{(s)}$/$p^{(s)}_{Z}$ are thresholds for photon loss, photon indistinguishability from different emitters, and spin depolarizing/$Z$ error probability, respectively.
@@
To suppress these errors, in Sec. VI we discuss the use of code concatenation, redundantly encoding each qubit of the linear cluster state and performing logical fusions between them. We consider two types of logical fusion, based on static repetition codes (REP) and repeat-until-success gates (RUS), and elucidate the relationship between the two. The thresholds of our construction to various noise models are then presented in Sec. VII, before suggesting directions for future work.
@@ IV Fault-tolerance under measurement errors
Simulations are performed by Monte-Carlo sampling of different configurations of fusion error and erasure. For each data point, $10^{4}$ samples are drawn to estimate the logical error rate, which is a fraction of the number of samples with logical errors or erasures. The lattice is simulated using periodic boundary conditions, as we are interested primarily in the behaviour of the bulk. To identify the region of fault rates that can be corrected by the architecture, we linearly parameterise $p_{\text{err}}$ and $p_{\text{eras}}$ by an error parameter $x$, and calculate logical error rates for various lattice sizes $L$ as $x$ is varied.
@@ VI Local repetition codes
Figure 6: Logical fusion protocols. (a) With a static repetition (REP) code size of $m=3$, a logical fusion equates to applying $m$ transversal physical fusions between encoded qubits of each resource state. (b) Repeat-until-success (RUS) logical fusion performs a minimal number of physical fusions to recover one or two logical outcomes. If the first physical fusion fails, another pair of photons is emitted for a second attempt. If one of the photons is lost, the next fusion will be biased in $ZZ$ [10] to guarantee an outcome in $\overline{ZZ}$.
VI.1 Logical Fusion with repetition codes
With a local encoding, the physical fusions are substituted for logical fusion measurements. For an $m$-qubit repetition code, the fusion group $\mathcal{F}$ associated with a logical fusion between two qubits $a$ and $b$, is
@@ VI.1 Logical Fusion with repetition codes
where $a_{j}$ refers to the $j$th physical qubit in the logical qubit $a$. These operators can be recovered by performing fusions transversally (Fig. 6a), such that each qubit from one code undergoes a physical fusion with its equivalent in the other.
@@
VI.2 Repeat-until-success logical fusions
It is well known that adaptivity in physical fusion measurements can greatly improve the recovery probability for logical fusion outcomes over static methods [16, 14, 13]. So far we have considered logical fusions with a fixed number of physical fusions $m$ defined by the repetition code, but a more resource-efficient approach is possible using a repeat-until-success (RUS) fusion. This can be thought of as a variable-size repetition code, where we only generate the minimal size of code required to achieve a logical fusion.
The central idea of RUS logical fusion is to terminate further fusion attempts when one of the following conditions is met: (1) both $\overline{XX}$ and $\overline{ZZ}$ outcomes are recovered from prior attempts, (2) only $\overline{ZZ}$ is recovered, or (3) a maximum number of attempts $N$ is reached. This reduces the risk of a photon loss erasing the $\overline{XX}$ outcome, and the smaller average code size $m$ reduces the impact of $Z$ errors. If a loss occurs in a prior fusion attempt, $\overline{XX}$ is no longer recoverable, and the failure basis of the next physical fusion is switched to $ZZ$ [10] to recover $\overline{ZZ}$ with high probability (Fig. 6b).
@@ VI.2 Repeat-until-success logical fusions
Figure 7: Loss and indistinguishability thresholds. (a) Photon loss threshold as a function of number of fusion attempts per logical fusion ($m$ for REP and $N$ for RUS). Red solid circle represents RUS with $N$ re-attempts when emitters are re-initialized on the previous layer of fusions. (b) Indistinguishability between photons from different emitters required for fault-tolerance, assuming no photon loss.
The probabilities for different logical fusion outcomes of a single RUS fusion under photon loss and errors can be analytically derived, see Supplementary Note II. Using these analytical expressions, logical fusion outcomes are sampled to obtain a syndrome graph for calculating fault-tolerant thresholds.
@@ VII Fault-tolerant Thresholds
The loss-tolerance thresholds are presented in Fig. 7a, for various code sizes $m$ for REP and maximal number of physical fusion attempts $N$ for RUS. The highest REP loss threshold obtained is $\ell=2.29\%$ for $m=5$. For RUS logical fusion, we observe that $\ell$ improves significantly as the number of maximally allowed fusion attempts $N$ increases, achieving $\ell=7.4\%$ with $N=10$. Using a strategy where additional $N$ re-attempts of RUS are allowed with both emitters re-initialized from the previous layer of logical fusions (Supplementary Note III), $\ell$ can reach up to $8\%$.
@@
In the absence of loss, the RUS distinguishability threshold continues to increase with $N$, but saturates at $V=95.9\%$ (Fig. 7b) when $N=10$, whereas REP can tolerate $V=92\%$ with $m=10$. The reason for the difference in thresholds is that photon distinguishability between emitters introduces a $X$ error in the rotated fusion gate assumed for these logical fusions, which is well-tolerated by the repetition code through majority voting. This strategy is not available using RUS fusions which are terminated once the first $ZZ$ outcome is obtained, thus leading to the difference in thresholds observed here.
@@
For thresholds under spin noise, simulations are performed in two steps: (1) the logical fusion outcomes are sampled without photon loss. With only fusion failure $P_{\text{fail}}=50\%$, RUS results in either (i) $\overline{XX}$, or (ii) both $\overline{XX}$ and $\overline{ZZ}$ outcomes. (2) For each logical fusion that ends up in case (i), all physical fusions failed thus $m=N$. For case (ii), the number of fusion attempts $m\leq N$ is sampled according to the probability $(1-P_{\text{fail}})P_{\text{fail}}^{m-1}$. A list of values of $m$ is then obtained and used to extract logical fusion errors based on a pre-sampled list of photon errors originated from spin noise during resource generation (Sec. V.2). It is important to note that the simulation only accounts for spin decoherence of the two emitters involved in the current logical fusion, whereas other emitters are assumed to be perfectly coherent during the RUS gate.
@@
Fig. 8a plots the simulated spin $Z$ error threshold with the number of fusion attempts per logical fusion, with $p^{(s)}_{X}=p^{(s)}_{Y}=0$. We observe that for spin errors occurred during the generation of encoded linear chains, the RUS threshold increases up to $p^{(s)}_{Z}=0.59\%$ at $N=10$ whereas for REP it peaks at $p^{(s)}_{Z}=0.19\%$ then steadily decreases to $0.095\%$. As discussed in Sec. V.2, a spin $Z$ error will localize as a $Z$ error in the next emitted photon. For REP the number of fusion attempts $m$ is fixed, such that when $m$ grows the probability of flipping $\overline{XX}$ also increases, and this leads to the decreasing threshold for high $m$. In the limit of low $m$, however, more fusion attempts increases the fusion success probability resulting in a more connected and hence more robust lattice, with an optimal threshold at $m=5$. For RUS, most logical fusions terminate in a few attempts assuming no photon loss (Supplementary Note II), so that only $Z$ errors occurring in the photons emitted during those attempts contribute to a flip in $\overline{XX}$. As a consequence, the spin $Z$ error threshold saturates as we increase $N$.
@@
For spin depolarizing error (Fig. 8b), both thresholds again increase for low number of fusion attempts thanks to a better connectivity of the graph as explained above. The RUS (REP) threshold peaks at $p^{(s)}=0.2\%$ ($0.125\%$) then gradually decreases. This can be understood by first considering the propagation of spin $X$/$Y$ errors during resource state generation into photon errors, before mapping these photon errors to logical fusion errors. A spin $X$/$Y$ error occurred while generating the first $k\leq m$ photons in a logical qubit will lead to $X/Y$ errors in the subsequently generated $m-k$ photons  [^1] More precisely, a spin $Y$ error results in a $Y$ error on the first emitted photon and $X$ errors on the next generated photons in the current logical qubit $i$, as well as a $Z$ error on the first photon of the next qubit $i+1$., as well as a $Z$ photon error on the first photon of the next logical qubit (Sec. V.2). This implies that the probability of flipping the $ZZ$ outcome (hence $\overline{ZZ}$) increases with the number of fusion attempts, since a larger $m$ results in a greater probability of $X/Y$ spin errors within an $m$-photon logical qubit. Both error thresholds therefore reduce at higher values of $m$/$N$. Moreover, for REP, majority voting becomes ineffective at correcting erroneous $ZZ$ outcome due to a high weight of $X/Y$ errors from a single spin error.
@@
Figure 8: Spin error thresholds. (a) Spin $Z$ error threshold $p^{(s)}_{Z}$ as a function of number of fusion attempts per logical fusion ($m$ for REP and $N$ for RUS), with $p^{(s)}_{X}=p^{(s)}_{Y}=0$. (b) Spin depolarizing error threshold $p^{(s)}$ where each spin Pauli error probability is identical: $p^{(s)}_{X}=p^{(s)}_{Y}=p^{(s)}_{Z}=p^{(s)}/3$.
@@ VIII Discussion
Photon loss still remains to be the main bottleneck for FBQC, but recent experimental breakthroughs in photonic hardware [35, 36] and further optimization in the architecture hold promise to bridge this gap. For instance, deterministic entanglement between two emitters is possible by attempting RUS fusions between their respective coupled emitters, followed by entanglement swapping upon fusion success. This would significantly improve the logical fusion success probability and hence its resistance to photon loss [37, 15].
Our analysis provides the foundations for capturing realistic errors during resource state generation in emitter-based FBQC architectures. Further study will explore more detailed noise models including hardware-specific error processes, such as imperfect Raman pulses [5], finite optical cyclicity [8] and non-Markovian spin noises relevant to photon emitters [38]. Our work therefore creates a clear path for further development of both quantum photonic hardware and fusion-based architectures based on deterministic quantum emitters.
@@ I Proof for edge and node split rules
with $Z_{\mathcal{N}_{1}\setminus\{2\}}$ ($Z_{\mathcal{N}_{7}\setminus\{6\}}$) in the first graph is identical to $Z_{\mathcal{N}_{1}\setminus\{4\}}$ ($Z_{\mathcal{N}_{7}\setminus\{8\}}$) in the final graph.
II Erasure and error rates of Logical fusions
Here we calculate the probability of obtaining or recovering logical outcomes $\overline{XX}$ and $\overline{ZZ}$ in a logical fusion, using REP and RUS in the presence of photon losses. The logical fusion error rates are also computed, with $X$ and $Z$ errors applied to each physical fusion. In general, with an end-to-end efficiency of $\eta$ and a fusion failure rate of $P_{\text{fail}}$, for a physical fusion there are three possible events:

### v3

@@ I Introduction
Here, we present a fusion-based architecture tailored to quantum emitters and analyze how relevant physical processes affect its fault-tolerant performance. The architecture is based on the foliated Floquet color code (FFCC) [11] where each resource state can be deterministically generated by a single quantum emitter, with high tolerance to photon loss and qubit errors. The fault-tolerant performance of FFCC was previously analyzed using a phenomenological error model for fusions, in which erasure and error occurs independently for each fusion measurement [11], without considering physical noises dominant in quantum emitters. We develop a framework to describe how these noises propagate to the logical qubit level of the architecture and evaluate its performance in terms of fault-tolerant thresholds. Specifically, we investigate photon distinguishability, Markovian spin noise in spin-photon interfaces, and photon loss. We show that our tailored architectures are highly resilient to these noises, as summarized in Table 1.
@@
Table 1: Summary of main results. sFFCC: synchronous foliated Floquet color code lattice. REP: encoded fusion with repetition code of fixed code size $m$. RUS: repeat-until-success encoded fusion with varying code size $m\leq N$ where $N$ is the maximum number of physical fusion attempts. $\ell$, $V$ and $p^{(s)}$/$p^{(s)}_{Z}$ are thresholds for photon loss, photon indistinguishability from different emitters, and spin depolarizing/$Z$ error probability, respectively.
@@
To suppress these errors, in Sec. VI we discuss the use of code concatenation, redundantly encoding each qubit of the linear cluster state and performing encoded fusions between them. We consider two types of encoded fusion, based on static repetition codes (REP) and repeat-until-success gates (RUS), and elucidate the relationship between the two. The thresholds of our construction to various noise models are then presented in Sec. VII, before suggesting directions for future work.
@@ IV Fault-tolerance under measurement errors
Simulations are performed by Monte-Carlo sampling of different configurations of fusion error and erasure. For each data point, $N_{\text{trials}}=10^{4}$ trials are used to estimate the logical error rate $R=N_{\text{fail}}/N_{\text{trials}}$, where $N_{\text{fail}}$ is the number of trials when the decoding fails due to either logical errors or erasures. The lattice is simulated using periodic boundary conditions, as we are interested primarily in the behaviour of the bulk. To identify the region of fault rates that can be corrected by the architecture, we linearly parameterise $p_{\text{err}}$ and $p_{\text{eras}}$ by an error parameter $x$, and calculate logical error rates for various lattice sizes $L$ as $x$ is varied.
@@ VI Local repetition codes
Figure 6: Encoded fusion protocols. (a) With a static repetition (REP) code size of $m=3$, an encoded fusion equates to applying $m$ transversal physical fusions between encoded qubits of each resource state. (b) Repeat-until-success (RUS) encoded fusion performs a minimal number of physical fusions to recover one or two measurement outcomes. If the first physical fusion fails, another pair of photons is emitted for a second attempt. If one of the photons is lost, the next fusion will be biased in $ZZ$ [10] to guarantee an outcome in $\overline{ZZ}$.
VI.1 Encoded Fusion with repetition codes
With a local encoding, the physical fusions are substituted for encoded fusion measurements. For an $m$-qubit repetition code, the fusion group $\mathcal{F}$ associated with an encoded fusion between two qubits $a$ and $b$, is
@@ VI.1 Encoded Fusion with repetition codes
where $a_{j}$ refers to the $j$th physical qubit in the encoded qubit $a$. These operators can be recovered by performing fusions transversally (Fig. 6a), such that each qubit from one code undergoes a physical fusion with its equivalent in the other.
@@
VI.2 Repeat-until-success encoded fusions
It is well known that adaptivity in physical fusion measurements can greatly improve the recovery probability for encoded fusion outcomes over static methods [16, 14, 13]. So far we have considered encoded fusions with a fixed number of physical fusions $m$ defined by the repetition code, but a more resource-efficient approach is possible using a repeat-until-success (RUS) fusion. This can be thought of as a variable-size repetition code, where we only generate the minimal size of code required to achieve an encoded fusion.
The central idea of RUS encoded fusion is to terminate further fusion attempts when one of the following conditions is met: (1) both $\overline{XX}$ and $\overline{ZZ}$ outcomes are recovered from prior attempts, (2) only $\overline{ZZ}$ is recovered, or (3) a maximum number of attempts $N$ is reached. This reduces the risk of a photon loss erasing the $\overline{XX}$ outcome, and the smaller average code size $m$ reduces the impact of $Z$ errors. If a loss occurs in a prior fusion attempt, $\overline{XX}$ is no longer recoverable, and the failure basis of the next physical fusion is switched to $ZZ$ [10] to recover $\overline{ZZ}$ with high probability (Fig. 6b).
@@ VI.2 Repeat-until-success encoded fusions
Figure 7: Loss and indistinguishability thresholds. (a) Photon loss threshold as a function of number of fusion attempts per encoded fusion ($m$ for REP and $N$ for RUS). Red solid circle represents RUS with $N$ re-attempts when emitters are re-initialized on the previous layer of fusions. (b) Indistinguishability between photons from different emitters required for fault-tolerance, assuming no photon loss.
The probabilities for different encoded fusion outcomes of a single RUS fusion under photon loss and errors can be analytically derived, see Supplementary Note II. Using these analytical expressions, encoded fusion outcomes are sampled to obtain a syndrome graph for calculating fault-tolerant thresholds.
@@ VII Fault-tolerant Thresholds
The loss-tolerance thresholds are presented in Fig. 7a, for various code sizes $m$ for REP and maximal number of physical fusion attempts $N$ for RUS. The highest REP loss threshold obtained is $\ell=2.29\%$ for $m=5$. For RUS encoded fusion, we observe that $\ell$ improves significantly as the number of maximally allowed fusion attempts $N$ increases, achieving $\ell=7.4\%$ with $N=10$. $\ell$ can be further improved to $8\%$, by using a strategy where additional $N$ re-attempts of RUS are allowed with both emitters re-initialized from the previous layer of encoded fusions. This trick stems from the observation that a biased physical fusion requires performing single-qubit $Z$ measurements on both photons to generate the $\overline{ZZ}$ outcome. Since the emitted photon and its entangled spin share the stabilizer $ZZ$, measuring the photon in $Z$ effectively projects the spin in $Z$-basis, which re-initializes it. If two emitters were re-initialized from the previous layer of encoded fusions, they are detached from any correlations and hence $N$ additional re-attempts of RUS can be applied to make the next encoded fusion near-deterministic (Supplementary Note III).
@@
In the absence of loss, the RUS distinguishability threshold continues to increase with $N$, but saturates at $V=95.9\%$ (Fig. 7b) when $N=10$, whereas REP can tolerate $V=92\%$ with $m=10$. The reason for the difference in thresholds is that photon distinguishability between emitters introduces a $X$ error in the rotated fusion gate assumed for these encoded fusions, which is well-tolerated by the repetition code through majority voting. This strategy is not available using RUS fusions which are terminated once the first $ZZ$ outcome is obtained, thus leading to the difference in thresholds observed here.
@@
For thresholds under spin noise, simulations are performed in two steps: (1) the encoded fusion outcomes are sampled without photon loss. With only fusion failure $P_{\text{fail}}=50\%$, RUS results in either (i) $\overline{XX}$, or (ii) both $\overline{XX}$ and $\overline{ZZ}$ outcomes. (2) For each encoded fusion that ends up in case (i), all physical fusions failed thus $m=N$. For case (ii), the number of fusion attempts $m\leq N$ is sampled according to the probability $(1-P_{\text{fail}})P_{\text{fail}}^{m-1}$. A list of values of $m$ is then obtained and used to extract encoded fusion errors based on a pre-sampled list of photon errors originated from spin noise during resource generation (Sec. V.2). It is important to note that the simulation only accounts for spin decoherence of the two emitters involved in the current encoded fusion, whereas other emitters are assumed to be perfectly coherent during the RUS gate.
@@
Fig. 8a plots the simulated spin $Z$ error threshold with the number of fusion attempts per encoded fusion, with $p^{(s)}_{X}=p^{(s)}_{Y}=0$. We observe that for spin errors occurred during the generation of encoded linear chains, the RUS threshold increases up to $p^{(s)}_{Z}=0.59\%$ at $N=10$ whereas for REP it peaks at $p^{(s)}_{Z}=0.19\%$ then steadily decreases to $0.095\%$. As discussed in Sec. V.2, a spin $Z$ error will localize as a $Z$ error in the next emitted photon. For REP the number of fusion attempts $m$ is fixed, such that when $m$ grows the probability of flipping $\overline{XX}$ also increases, and this leads to the decreasing threshold for high $m$. In the limit of low $m$, however, more fusion attempts increases the fusion success probability resulting in a more connected and hence more robust lattice, with an optimal threshold at $m=5$. For RUS, most encoded fusions terminate in a few attempts assuming no photon loss (Supplementary Note II), so that only $Z$ errors occurring in the photons emitted during those attempts contribute to a flip in $\overline{XX}$. As a consequence, the spin $Z$ error threshold saturates as we increase $N$.
@@
For spin depolarizing error (Fig. 8b), both thresholds again increase for low number of fusion attempts thanks to a better connectivity of the graph as explained above. The RUS (REP) threshold peaks at $p^{(s)}=0.2\%$ ($0.125\%$) then gradually decreases. This can be understood by first considering the propagation of spin $X$/$Y$ errors during resource state generation into photon errors, before mapping these photon errors to encoded fusion errors. A spin $X$/$Y$ error occurred while generating the first $k\leq m$ photons in an encoded qubit will lead to $X/Y$ errors in the subsequently generated $m-k$ photons  [^1] More precisely, a spin $Y$ error results in a $Y$ error on the first emitted photon and $X$ errors on the next generated photons in the current encoded qubit $i$, as well as a $Z$ error on the first photon of the next qubit $i+1$., as well as a $Z$ photon error on the first photon of the next encoded qubit (Sec. V.2). This implies that the probability of flipping the $ZZ$ outcome (hence $\overline{ZZ}$) increases with the number of fusion attempts, since a larger $m$ results in a greater probability of $X/Y$ spin errors within an $m$-photon encoded qubit. Both error thresholds therefore reduce at higher values of $m$/$N$. Moreover, for REP, majority voting becomes ineffective at correcting erroneous $ZZ$ outcome due to a high weight of $X/Y$ errors from a single spin error.
In practice, the above considered errors will likely be present simultaneously during computation. In Supplementary Note IV, we have considered a combination of photon loss, distinguishability and spin depolarizing errors, and simulated the corresponding fault-tolerant region using RUS fusions.
@@
Figure 8: Spin error thresholds. (a) Spin $Z$ error threshold $p^{(s)}_{Z}$ as a function of number of fusion attempts per encoded fusion ($m$ for REP and $N$ for RUS), with $p^{(s)}_{X}=p^{(s)}_{Y}=0$. (b) Spin depolarizing error threshold $p^{(s)}$ where each spin Pauli error probability is identical: $p^{(s)}_{X}=p^{(s)}_{Y}=p^{(s)}_{Z}=p^{(s)}/3$.
@@ VIII Discussion
Photon loss still remains to be the primary bottleneck for FBQC, with the highest demonstrated outcoupling efficiency currently limited to $71\%$ [35]. However, recent experimental breakthroughs in photonic hardware [36, 37] and further architectural optimizations hold promise to bridge this gap. For instance, deterministic entanglement between two emitters is possible by attempting RUS fusions between their respective coupled emitters, followed by entanglement swapping upon fusion success. This would significantly improve the encoded fusion success probability and hence its resistance to photon loss [38, 15]. A detailed comparison between the aforementioned thresholds and current experimental capabilities is provided in Supplementary Note V.
Our analysis provides the foundations for capturing realistic errors during resource state generation in emitter-based FBQC architectures. Further study will explore more detailed noise models including hardware-specific error processes, such as imperfect Raman pulses [5], finite optical cyclicity [8] and non-Markovian spin noises relevant to photon emitters [39].
While architectures with RUS fusions demonstrated greater tolerance to loss and spin errors than that with REP (static repetition code) fusions, they may also be more experimentally demanding. For RUS fusions between two resource state generators, no fiber delay is required, but the failure basis needs to be adaptively switched based on the previous physical fusion outcome. In the REP architecture, fusions on the same layer can in principle be performed synchronously with less resource state generators through interleaving [40]. However, this requires additional fiber delays which, in combination with the lower loss threshold ($\ell=2.3\%$) of the REP architecture, can pose significant challenges.
Despite these experimental hurdles, the synchronous FFCC architecture requires only a fixed number of optical switches per resource state generator, independent of the lattice size. This is in contrast to the original FFCC architecture [11], which requires an increasing number of switches for fusions between different resource states at larger lattice sizes, as the resource states percolate across multiple unit cells. The practical implementation of these architectures will be a key focus for future research directions. Our work, therefore, creates a clear path for further development of both quantum photonic hardware and fusion-based architectures leveraging deterministic quantum emitters.
@@ Acknowledgement
Data Availability
The data that support the findings of this article are openly available [42].
@@ I Proof for edge and node split rules
where $Z_{\mathcal{N}_{1}\setminus\{2\}}$ ($Z_{\mathcal{N}_{7}\setminus\{6\}}$) in the first graph is identical to $Z_{\mathcal{N}_{1}\setminus\{4\}}$ ($Z_{\mathcal{N}_{7}\setminus\{8\}}$) in the final graph.
II Erasure and error rates of encoded fusions
Here we calculate the probability of obtaining or recovering measurement outcomes $\overline{XX}$ and $\overline{ZZ}$ in an encoded fusion, using REP and RUS in the presence of photon loss. The encoded fusion error rates are also computed, with $X$ and $Z$ errors applied to each physical fusion. In general, with an end-to-end efficiency of $\eta$ and a fusion failure rate of $P_{\text{fail}}$, for a physical fusion there are three possible events: